OpenStudy (anonymous):

Nick is now seven times as old as Harry. In 6 years, Nick will be five times as old as Harry. How old will Harry be in 6 years? 12 16 18

OpenStudy (anonymous):

18

OpenStudy (apoorvk):

Let Nick's present age be 'x', and Harry's be 'y' so, since 'x' is seven times that of 'y', x=7y ....{i} After 6 years, Nick's age = x+6 = 7y+6 ...{from 'i'} Harry's age after 6 years = y + 6 now, it's been stated that after 6 years, Nick shall be 5 times as old as Harry. So, (7y+6) = 5(y+6) Solve this^!

OpenStudy (anonymous):

really!!

OpenStudy (lgbasallote):

Yes!!!

OpenStudy (anonymous):

combine like terms first?

OpenStudy (lgbasallote):

Distribute 5 first

OpenStudy (apoorvk):

yes yes!

OpenStudy (apoorvk):

and distribute '5' first, as Monsieur Basallote says!

OpenStudy (lgbasallote):

Combine second

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

multiply 5 by 6 to right

OpenStudy (anonymous):

7y+6=5y+30 2y =24 y=12

OpenStudy (apoorvk):

yes multiply 5 by 6 to right and add a y multiplied by 5 to it as well.

OpenStudy (anonymous):

well this is now pointless because i already have the answer

OpenStudy (apoorvk):

Hmm maybe. But did you understand it? Or rather you feel you could have done that yourself?

OpenStudy (anonymous):

i actually could have done that one i think